Assume that a bank receives a deposit of $1,000 in cash, puts aside $200 as required reserves, and makes a loan of $800, these transactions imply that:

the money multiplier is 10.

the money multiplier is 5.

the money multiplier is 4.

the money multiplier is 2.

the money multiplier is 2.
Answer #1

Reserve ratio = reserve maintained / deposit = 200/1000 =0.2 = r

Money multiplier = 1/reserve ratio = 1/r = 1/0.2 = 5

Ans 5

If any doubt ask